# StreamBench: Towards Benchmarking Continuous Improvement of Language Agents

### v1

@@ Abstract
Recent works have shown that large language model (LLM) agents are able to improve themselves from experience, which is an important ability for continuous enhancement post-deployment. However, existing benchmarks primarily evaluate their innate capabilities and do not assess their ability to improve over time. To address this gap, we introduce StreamBench, a pioneering benchmark designed to evaluate the continuous improvement of LLM agents over an input-feedback sequence. StreamBench simulates an online learning environment where LLMs receive a continuous flow of feedback stream and iteratively enhance their performance. In addition, we propose several simple yet effective baselines for improving LLMs on StreamBench, and provide a comprehensive analysis to identify critical components that contribute to successful streaming strategies. Our work serves as a stepping stone towards developing effective online learning strategies for LLMs, paving the way for more adaptive AI systems in streaming scenarios.
@@ 3.1 General setup
Most public datasets are inherently static, meaning each instance does not have a time-related dependency. To adapt them for our streaming setup, we serialize each selected dataset in Section 3.2 by assigning a time step to each instance. To avoid arbitrary sequence assignment in the original datasets, we randomly shuffle each dataset using a fixed random seed. We release each dataset’s assigned sequence obtained by this random seed in the supplementary materials to ensure reproducibility on StreamBench. Additionally, to ensure the robustness of our evaluation, we conduct experiments on different shuffled sequences with five random seeds, as discussed in Section 5.2.
@@ 5.3 Would stronger LLMs still benefit from streaming?
To evaluate whether stronger LLMs still benefit from streaming, we tested two latest models from the GPT and Gemini families: gpt-4o-2024-05-13 and gemini-1.5-flash-preview-0514. Due to the high cost, we only run two baselines shown in Table 3. Despite their extraordinary zero-shot capabilities, we found that with Self-StreamICL, these stronger models still showed significant performance improvements. This demonstrates that even the most advanced models can leverage the information from streaming data to further enhance their performance across diverse tasks.
Table 3: Performance of latest models on StreamBench.
@@
Zero-Shot 69.31 48.96 33.05 49.70 69.60 54.08 59.33
Self-StreamICL 78.02 56.31 40.87 52.30 76.53 81.01 64.07
gpt-4o-2024-05-13
Zero-Shot 74.99 54.72 33.70 52.70 72.80 68.14 65.07
Self-StreamICL 81.14 58.39 43.09 60.90 76.53 93.14 66.80
@@ 5.4 Cost analysis
The input and output token usage of the all LLMs and baselines is listed in Appendix E for future benchmark users to estimate the cost when running StreamBench.
@@ Appendix D Detailed ablation study results
Figure 8: Correctness ablations of gpt-3.5-turbo-0125. The y-axis denotes performance difference from zero-shot.
Figure 9: Correctness ablations of gemini-1.0-pro-001. The y-axis denotes performance difference from zero-shot.
Figure 10: Correctness ablations of claude-3-haiku-20240307. The y-axis denotes performance difference from zero-shot.
@@ Appendix E Token cost breakdown for each LLM endpoint
Table 8, 9, 10, and 11 shows how many millions of input, output tokens are used by gpt-3.5-turbo-0125, gemini-1.0-pro-001, claude-3-haiku-20240307, and the latest models. The cost of MAM-StreamICL is simply the averaged cost of the three LLMs due to the round-robin algorithm. As most LLM endpoints use input and output tokens to charge usage fees, we provide this information for benchmark users to estimate the cost for running StreamBench.
@@ Appendix F Supplementary materials
The content in this section is provided as a separate PDF, as specified in the submission instructions. The supplementary materials include details such as preprocessing of datasets, prompts of each baseline method, and code to reproduce the experiments.

### v2

@@ Abstract
Recent works have shown that large language model (LLM) agents are able to improve themselves from experience, which is an important ability for continuous enhancement post-deployment. However, existing benchmarks primarily evaluate their innate capabilities and do not assess their ability to improve over time. To address this gap, we introduce StreamBench, a pioneering benchmark designed to evaluate the continuous improvement of LLM agents over an input-feedback sequence. StreamBench simulates an online learning environment where LLMs receive a continuous flow of feedback stream and iteratively enhance their performance. In addition, we propose several simple yet effective baselines for improving LLMs on StreamBench, and provide a comprehensive analysis to identify critical components that contribute to successful streaming strategies. Our work serves as a stepping stone towards developing effective online learning strategies for LLMs, paving the way for more adaptive AI systems in streaming scenarios. Source code: https://github.com/stream-bench/stream-bench. Benchmark website: https://stream-bench.github.io.
@@ 3.1 General setup
Most public datasets are inherently static, meaning each instance does not have a time-related dependency. To adapt them for our streaming setup, we serialize each selected dataset in Section 3.2 by assigning a time step to each instance. To avoid arbitrary sequence assignment in the original datasets, we randomly shuffle each dataset using a fixed random seed. We release each dataset’s assigned sequence obtained by this random seed in the supplementary materials to ensure reproducibility on StreamBench. Additionally, to ensure the robustness of our evaluation, we conduct experiments on different shuffled sequences with five random seeds, as discussed in Section 5.2. We also discuss the effects of distributional shifts in Appendix C.2.
@@ 5.3 Would stronger LLMs still benefit from streaming?
To evaluate whether stronger LLMs still benefit from streaming, we tested two newer models: gpt-4o-2024-08-06 and gemini-1.5-flash-001. Due to the high cost, we only run the methods shown in Table 3. We found that with Self-StreamICL, these stronger models still showed significant performance improvements. This demonstrates that even the most advanced models can leverage the information from streaming data to further enhance their performance across diverse tasks.
Table 3: Performance of gpt-4o-2024-08-06 and gemini-1.5-flash-001 on StreamBench.
@@
Zero-shot 69.63 48.26 33.83 50.20 69.47 58.90 60.60
Few-shot 71.40 49.35 37.03 50.60 72.13 73.58 64.87
CoT 72.52 52.73 35.14 44.80 68.00 64.06 63.13
Self-StreamICL 77.83 56.21 41.20 52.20 75.07 86.34 65.20
gpt-4o-2024-08-06
Zero-shot 73.54 53.33 34.42 54.90 72.40 70.64 65.53
Few-shot 76.85 57.60 36.25 52.30 71.47 83.45 66.87
CoT 72.52 54.82 31.16 41.90 66.80 73.02 62.80
Self-StreamICL 80.58 59.19 42.63 59.40 76.27 92.01 67.00
@@ 5.4 Cost analysis
For benchmark users to estimate the cost, the token usage of all baselines is listed in Appendix E.
@@ 8 Limitations
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See the experiment results in Section 4.3.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 8.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Please refer to the supplementary materials.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Please refer to Appendix A and  F.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See the results in Figure 4.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 5.4.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See Section 3.2.
    2. (b)
      Did you mention the license of the assets? [Yes] See the supplementary materials.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] We include code in the supplementary materials.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See the supplementary materials.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ Appendix A Hyperparameters
We also analyze how different text embeddings used in memory correlate with streaming performance in Table 4. We observe that within the same text encoder family (bge), the larger model (109M parameters) generally delivers better performance. However, smaller models (22.7M parameters) can also achieve strong results, indicating that each LLM may benefit from a specific encoder.
Table 4: Performance of Self-StreamICL (implemented with different text encoders and LLMs) on the DDXPlus dataset.
Text encoder / LLMs gpt-3.5-turbo-0125 claude-3-haiku gemini-1.5-flash-001
all-MiniLM-L6-v2 (22.7M) 63.61 78.91 83.50
bge-small-en-v1.5 (33.4M) 63.55 75.51 83.90
bge-base-en-v1.5 (109M) 66.16 76.02 86.34
@@ Appendix C Robustness to different streaming sequences
C.1 Performance on sequences shuffled by five different random seeds
@@
C.2 Performance on the sequence with distributional shifts
Since we randomly assign time steps to each instance, our main results in Table 2 simulates scenarios where each instance in the streaming sequence is drawn from the same distribution. However, it is important to investigate how well streaming methods perform on sequences with distributional shifts. To this end, we conduct experiments on BIRD [19] by arranging instances from the same database consecutively (DB1, DB2, …, DB11), resulting in 10 distributional shifts during streaming. The results are shown in Table 8. The key finding is that Self-StreamICL still outperforms non-streaming baselines, and the method’s performance does not differ drastically when distributional shifts occur.
Table 8: Performance of Self-StreamICL with different LLMs on BIRD.
Method / LLM gpt-3.5 gemini-1.0-pro claude-3-haiku gemini-1.5-flash gpt-4o
Zero-Shot 29.75 28.16 30.90 33.83 34.42
Few-Shot 28.94 31.49 30.77 37.03 36.25
CoT 29.21 21.71 30.77 35.14 31.16
Self-StreamICL (no shifts) 35.07 33.25 37.61 41.20 42.63
Self-StreamICL (w/ shifts) 36.31 32.60 36.57 40.48 43.09
@@ Appendix D Detailed ablation study results
Figure 8: Correctness ablations of gpt-3.5-turbo-0125.
Figure 9: Correctness ablations of gemini-1.0-pro-001.
Figure 10: Correctness ablations of claude-3-haiku-20240307.
@@ Appendix E Token cost breakdown for each LLM endpoint
Table 10, 11, 12, and 13 shows how many millions of input, output tokens are used by gpt-3.5-turbo-0125, gemini-1.0-pro-001, claude-3-haiku-20240307, and the latest models. The cost of MAM-StreamICL is simply the averaged cost of the three LLMs due to the round-robin algorithm. As most LLM endpoints use input and output tokens to charge usage fees, we provide this information for benchmark users to estimate the cost for running StreamBench.
@@ Appendix F Supplementary materials
The supplementary materials include details such as preprocessing of datasets, prompts of each baseline method, and code to reproduce the experiments.